Minnie is holding a yardstick perpendicular to the ground. The shadow of the yardstick is 27 inches long. Minnie’s shadow is 4 feet 6 inches long. How tall is Minnie?
Answer by stanbon(75887)

Minnie is holding a yardstick perpendicular to the ground. The shadow of the yardstick is 27 inches long. Minnie’s shadow is 4 feet 6 inches long. How tall is Minnie?
------------
Note: 4' 6" = 54"
Proportion:
x/54 = 36/27
x = 54(36/27)
x = 2*36
x = 72"
Minnie is 6 ft tall
=========================
Cheers,
Stan H.
===========